# On $\mathcal{F}$-semi-transitive families of operators

### v4

@@ 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [25]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}^{(\ell)}$, i.e.
@@
We will now present an application of Theorem 4.2 in this concrete case.
@@
Under the above notation and assumptions, the following statements are equivalent:
@@
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators
$\{D_{j}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{j,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{j,N}^{(k)}\}_{k\in\mathbb{N}}.$
in $\mathcal{C}$ for every $j\in\mathbb{Z}$ with $J\leq j\leq J$ such that
@@
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0.$
@@
In order to construct concrete examples satisfying the conditions $(2)$ of Corollary 4.3, we need first the following technical result.
@@
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0.,$
then he sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
@@
The main idea in this proof is motivated by the proofs of [25, Theorem 3.2] and [26, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
@@
By some calculations and arguments similar to those in the proof of [25, Theorem 3.2] and [26, Proposition 2.7], we can conclude that the sequences of operators
@@
Now we are ready to give an example where the conditions of Corollary 4.3 and Proposition 4.4 are satisfied.

### v5

@@ 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
for all $n\in F$ and $\ell\in\{1,\dots,N\}$. ∎
Next, let $\{S_{n}\}_{n\in\mathbb{N}}$ be a sequence of bounded invertible linear operators on $H$ and $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be families of generalized weighted bilateral shifts on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section. For each $n\in F$ and $s\in\{1,\dots,N\}$ we let $S_{n}T_{n,s}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
$\big(S_{n}T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big):=\{S_{n}W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
It is not hard to check that the operator $S_{n}T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{C}).$ In fact, $S_{n}T_{n,s}$ is again a generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ with respective weights $\{S_{n}W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ for every $n\in F$ and $s\in\{1,\dots,N\}.$
We will say that $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-C*-semi-transitive , or shortly $d\mathcal{F}$-C*-semi-transitive , if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $n\in F$ there exists some invertible bounded linear operator $S_{n}$ on $H$ satisfying that
$\mathcal{O}\cap(S_{n}T_{n,1})^{-1}(V_{1})\cap\dots\cap(S_{n}T_{n,N})^{-1}(V_{N})\neq\varnothing.$
The next theorem can easily be deduced from Theorem 4.2.
Theorem 4.3.
(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
$\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|<\varepsilon,\text{ }\|W_{n,j}^{(s)\,-1}S_{n}^{-1}G_{n,j}^{(s)}\|<\varepsilon,$
$\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
Proof.
For the proof of $(1)\Rightarrow(2),$ we can proceed in exactly the same way as in the proof of Theorem 4.2 part $(1)\Rightarrow(2)$ by considering $S_{n}$ and $S_{n}^{-1}$ instead of $\lambda_{n}$ and $\tfrac{1}{\lambda_{n}},$ respectively.
For the proof of $(2)\Rightarrow(1),$ we can also proceed similarly as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$ However, the only difference is that in this case we let now for each $n\in F$, $v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else,}\end{cases}$
which clearly gives that
$\|(v_{n}^{(\ell)})\|\leq\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\|\|y^{(\ell)}\|.$
Then it suffices to also notice in this case that for each $n\in F$ and $\ell\in\{1,\dots,N\}$ it holds that
$\|S_{n}T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
(where $u_{n}$ is the same as in the proof of Theorem 4.2 in the part $(2)\Rightarrow(1)$) and and moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and $n\in F$ we have
$\|S_{n}T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
Finally, by the construction of $v_{n}^{(\ell)}$, it follows that
$S_{n}T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else,}$
for each $n\in F$ and $\ell\in\{1,\dots,N\}.$
For each $n\in F$, set
$x_{n}:=u_{n}+\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
$x_{n}\in\mathcal{O}\quad\text{and}\quad S_{n}T_{n,\ell}(x_{n})\in V_{\ell}$
@@
Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [26]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell}$, i.e.
@@
We will now present an application of Theorem 4.2 and Theorem 4.3 in this concrete case.
@@
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators in $\mathcal{C}$
$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
$\vdots$
$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
such that
@@
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0$
@@
Corollary 4.5.
(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N},$ a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ and sequences of operators in $\mathcal{C}$
$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
$\vdots$
$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
such that
and
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0,$
$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0,$
for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$ and satisfying in addition that for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0.$
In order to construct concrete examples satisfying the conditions in $(2)$ of Corollary 4.4 and Corollary 4.5, we need first the following technical results.
@@
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0,$
then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
@@
The main idea in this proof is motivated by the proofs of [26, Theorem 3.2] and [27, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
@@
By some calculations and arguments similar to those in the proof of [26, Theorem 3.2] and [27, Proposition 2.7], we can conclude that the sequences of operators
@@
Similarly, based on Theorem 4.3 and Corollary 4.5, we can prove the following proposition.
Proposition 4.7.
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H,$ a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}\,y_{s}^{(j)}\big\|=0,$
$\lim_{k\to\infty}\big\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}y_{s}^{(j)}\|=0,$
then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
Thanks to Proposition 4.6 we will now give an example where we will show that it can happen that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}$ and that at a same time the sequence of operators $\{\tilde{T}^{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$-semi-transitive, which illustrates the important difference between $\mathcal{F}$-semi-transitivity (and thus supercyclicity) of generalized bilateral weighted shifts on Hilbert C*-modules and supercyclicity of the classical bilateral weighted shifts on the sequence space $\ell^{2}.$